OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. R(4, 0), m = 5

OpenStudy (shamim):

the formula for a line is\[y-y _{1}=m(x-x _{1})\]

OpenStudy (shamim):

here given slope \[m=5,x _{1}=4,y _{1}=0\]

OpenStudy (shamim):

\[y-0=5(x-4)\]

OpenStudy (shamim):

try to solve it

OpenStudy (anonymous):

ah okay thanks

OpenStudy (shamim):

welcome

OpenStudy (anonymous):

so y=0+5x-20 right?